What is the value of Z in the solution of the system of equations?
2x + 3y + z = 9
x - 2y - z = 4
x - 3y + 2z = -3

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

What is the value of Z in the solution of the system of equations?
2x + 3y + z = 9
x - 2y - z = 4
x - 3y + 2z = -3
======================
I used the matrix function on my TI calculator
to get:
x = 4
y = 1
z = -2
=============
Cheers,
Stan H.

Answer by richwmiller(17219)

There is no Z in the system of equations. Z is not the same as z.